# Entanglement production through a cosmological bounce

### v1

@@ Abstract
In quantum gravity, it is expected that the Big Bang singularity is resolved and the universe undergoes a bounce. We find that for generic initial data, matter-gravity entanglement entropy rises rapidly during the bounce, declines, and then approaches a steady-state value following the bounce. These observations suggest that matter-gravity entanglement is a feature of the macroscopic universe and that there is no Second Law of entanglement entropy.
@@
Classical gravity is described by coupled equations for the degrees of freedom of gravity and matter. In the conventional application of quantization to gravity, the Hilbert space of states therefore takes the form of a bipartite system—the tensor product of gravity and matter Hilbert spaces: ${\cal H}_{QG}={\cal H}_{G}\otimes{\cal H}_{M}$. The matter Hilbert space is in general also a tensor product of the spaces for various species of matter. This tensor product structure could lead to matter-matter or matter-gravity entanglement. Either is a potential observable for quantum gravity [1, 2, 3], one which may play a role in understanding the emergence of quantum fields on curved spacetime from a theory of quantum gravity.
@@
where $t$ is the time in the dust gauge $T=t$ [12], and $\mu=1/4$ and $\mu=0$ realize two representations of the Hamiltonian operator used respectively in [11] and [10]. This Hamiltonian has a one-parameter family of self-adjoint extensions, all of which require the condition $\lim_{v\rightarrow 0}\psi(t,v,\varphi)=0$.
The Schrodinger equation (2) is readily solved by separating variables and assuming that the field $\varphi$ takes values in a circle of radius $R$; this is a convenient device since $R$ may be taken to be as large as desired. Solutions are then given by the expansion
@@
Each “mode” $n$ in the expansion (3) gives rise to a corresponding one-dimensional Schrodinger equation (4) with potential $V_{n}\equiv\alpha_{n}/v^{2}$. Thus, initial data $f(v)$ for (4) evolved with the potential $V_{n}$ gives a solution that may be represented as the product $\xi^{f}_{n}(t,v)\exp(i2n\pi\varphi/R)$.
To study entanglement dynamics with respect to the dust time $t$, we consider initial product states at $t=0$ of the form
$\psi(0,v,\varphi)=f(v)g(\varphi)=f(v)\left[\frac{1}{\sqrt{R}}\sum_{n=-\infty}^{\infty}c_{n}\exp\left(i\frac{2n\pi\varphi}{R}\right)\right]$ (5)
with $f(v)$ peaked at some initial volume $v_{0}$ with momentum $p_{0}$:
@@
For $p_{0}$ negative, the universe is initially hurtling toward the would-be singularity. The reduced density matrix for the evolution of such initial states, traced over the scalar field $\varphi$, is
@@
Our main result is the evolution of entanglement entropy through the bounce: its rapid rise through the bounce quantitatively captures the intuition of strong coupling between matter and gravity in regions of high curvature, and its subsequent relaxation to a constant nonzero value (noted earlier in [10]) indicates that matter-gravity entanglement may exist in our universe at present. These features are a consequence of the $p^{2}_{\varphi}/v^{2}$ coupling in the Hamiltonian (1), which indicates large coupling at small volume, and hence large entanglement. The qualitative features of the graphs in Fig. 2 are similar for other values of $(m,n)$, and the other parameters, indicating the general robustness of the entropy rise and decline through the bounce and the subsequent saturation.
This result, therefore, rules out the possibility of a Second Law for entanglement entropy, at least for the type of gravitational systems we study here. Indeed, a recent theorem arising from considerations of a general form of entanglement manipulation of quantum states also yields the result that there is no Second Law of entanglement entropy [14]; our calculation may be interpreted as an explicit form of entanglement manipulation coming from the cosmological Hamiltonian (1).

### v2

@@ Abstract
In quantum cosmology, it is expected that the Big Bang singularity is resolved and the universe undergoes a bounce. We find that for Gaussian initial states, matter-gravity entanglement entropy rises rapidly during the bounce, declines, and then approaches a steady-state value following the bounce. These observations suggest that matter-gravity entanglement is a feature of the macroscopic universe and that there is no Second Law of entanglement entropy.
@@
Classical gravity is described by coupled equations for the degrees of freedom of gravity and matter. In the conventional application of quantization to gravity, the Hilbert space of states therefore takes the form of a bipartite system—the tensor product of gravity and matter Hilbert spaces: ${\cal H}_{QG}={\cal H}_{G}\otimes{\cal H}_{M}$. The matter Hilbert space is in general also a tensor product of the spaces for various species of matter. This tensor product structure could lead to matter-matter or matter-gravity entanglement. Either is a potential observable for quantum gravity [1, 2, 3, 4], one which may play a role in understanding the emergence of quantum fields on curved spacetime from a theory of quantum gravity.
@@
where $t$ is the time in the dust gauge $T=t$ [13], and $\mu=1/4$ and $\mu=0$ realize two representations of the Hamiltonian operator used respectively in [12] and [11]. The Schrodinger equation (2) may be solved numerically as in [11] and also analytically by separation of variables as in [12] in terms of Bessel functions. In this work, we find it convenient to solve the equation numerically. To do so, we need to specify boundary conditions on the wave function. In order to calculate entanglement entropy through the bounce, it is useful to assume periodic boundary condition in the field direction $\varphi$; that is,
$\psi(t,v,\varphi+R)=\psi(t,v,\varphi),$ (3)
where $R$ is the period (which can be taken to be as large as desired). In order to have a well-defined numerical scheme, we also assume Dirichlet boundary conditions in the $v$ direction:
$\lim_{v\rightarrow 0}\psi(t,v,\varphi)=\lim_{v\rightarrow\infty}\psi(t,v,\varphi)=0.$ (4)
These boundary conditions ensure that the inner product between two solutions $\psi_{1}$ and $\psi_{2}$ of the Schrodinger equation is conserved:
[…]
Due to the periodicity in the field direction, solutions of the Schrodinger equation could be expressed as
@@
Each “mode” $n$ in the expansion (6) gives rise to a corresponding one-dimensional Schrodinger equation (7) with potential $V_{n}\equiv\alpha_{n}/v^{2}$. In order to calculate the wave function $\psi$ and the entanglement entropy, we solve equations of the form (7) numerically with boundary conditions [^1] It is interesting to note that reduced Schrödinger equation (7) can be solved numerically in terms of Bessel functions after introducing a mode decomposition of the form $\xi_{n}(t,v)=\sum_{j}e^{-iE_{nj}t}\zeta_{nj}(v)$. As discussed in detail in [14], it turns out that the boundary condition $\lim_{v\to 0}\xi_{n}=0$ is not enough to uniquely fix the mode solutions $\zeta_{nj}(v)$, something which results in a one-parameter family of self-adjoint extensions of the Hamiltonian. However, this ambiguity plays no role in the numeric solution of (7) since all self-adjoint extensions have $\lim_{v\to 0}\xi_{n}=0$. For a more comprehensive discussion of how various self-adjoint extensions modify the analytic mode solution of (2) in the context of a mode expansion, interested readers are directed to [12].
$\lim_{v\rightarrow 0}\xi_{n}(t,v)=\lim_{v\rightarrow\infty}\xi_{n}(t,v)=0.$ (8)
To study entanglement dynamics with respect to the dust time $t$, we consider initial product states for the total wavefunction at $t=0$ of the form
$\psi(0,v,\varphi)=f(v)g(\varphi)=f(v)\left[\frac{1}{\sqrt{R}}\sum_{n=-\infty}^{\infty}c_{n}\exp\left(i\frac{2n\pi\varphi}{R}\right)\right].$ (9)
This means we are interested in solutions of the reduced Schrodinger equation (7) with initial data $\xi_{n}(0,v)=f(v)$; we refer to solutions of (7) with this initial data as $\xi_{n}^{f}(t,v)$. We select $f(v)$ to be peaked at some initial volume $v_{0}$ with momentum $p_{0}$:
@@
The width $\sigma$ of these initial states is chosen small enough to ensure that the boundary condition $\lim_{v\to 0}\psi=0$ holds to within numerical accuracy at the initial time. For $p_{0}$ negative, the universe is initially hurtling toward the would-be singularity. The reduced density matrix for the evolution of such initial states, traced over the scalar field $\varphi$, is
@@
Our main result is the evolution of entanglement entropy through the bounce: its rapid rise through the bounce quantitatively captures the intuition of strong coupling between matter and gravity in regions of high curvature, and its subsequent relaxation to a constant nonzero value (noted earlier in [11]) indicates that matter-gravity entanglement may exist in our universe at present. These features are a consequence of the $p^{2}_{\varphi}/v^{2}$ coupling in the Hamiltonian (1), which indicates large coupling at small volume, and hence large entanglement. The qualitative features of the graphs in Fig. 2 are similar for other values of $(m,n)$, and the other parameters, indicating the general robustness of the entropy rise and decline through the bounce and the subsequent saturation. It is also noteworthy that these features of our results do not depend on the choice of self-adjoint extension of the Hamiltonian. Rather, they follow directly from the form of the Hamiltonian. This is because any initial product state must exhibit a rise in entanglement entropy from zero due to the increase in subsystem interaction as the volume decreases, followed by a freezing of the entropy as the volume increases after the bounce.
There is discussion in the literature of the possibility of a “Second Law for entanglement entropy” despite the many differences from the familiar thermodynamics version of the Second Law. A recent theorem arising from considerations of a general form of entanglement manipulation of quantum states yields the result that there is no Second Law of entanglement entropy [16]. In this context, our calculation supports this theorem, and may be interpreted as an explicit form of entanglement manipulation coming from the cosmological Hamiltonian (1); gravity and matter subsystems are less interacting due to the $1/v^{2}$ “coupling” at large volume, and more interacting at small volume.